OpenStudy (samigupta8):

the value of sin ^2 2π/3+cos ^2 5π/6-tan^ 3π/4 is

OpenStudy (anonymous):

2pi/3= pi/2+pi/6 5pi/6=pi-pi/6 3pi/4= pi-pi/4

OpenStudy (samigupta8):

ryt

OpenStudy (anonymous):

proceed now.

OpenStudy (samigupta8):

but it is coming as 1

OpenStudy (samigupta8):

and the answer as provided is 1/2

OpenStudy (anonymous):

its 1/2

OpenStudy (anonymous):